The electric potential at some point in space due to a point charge is 83 volts. If the charge is kept in the same place, what is the electric potential at a point which is 6 times further away compared to the first point? Answer in volts.
Answer #1

Electric potential depends inversely on the distance of observation point from the source charge hence if distance is made 6 times, potential will reduce by a factor of 1/6, so new potential is V = 83/6 = 13.83 Volts.